OpenStudy (anonymous):

Simplify the following expression using order of operations"PEMDAS" 4 + (6 × 4)20 - 32

OpenStudy (callisto):

4 + (6 × 4)20 - 32 = 4 + (24) 20 -32 = 4 + 480 -32 = ..?